OpenStudy (anonymous):

Examine the function f(x)=x+(3/x). Find the point on the curve at which the tangent lines pass through the point (1, 1).

hartnn (hartnn):

@byerskm2 Hi, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) Do you know how to find the slope of tangent to any curve f(x) ?

OpenStudy (anonymous):

Yes, I know the derivative is 1- 3/x^2 and the tangent line touching (1,1) is -2x+3 but then I got stuck

OpenStudy (anonymous):

How do I find the point on the the function line where the tangent line touches?

hartnn (hartnn):

u need to solve the equation of curve y=x+3/x and y=-2x+3 simultaneously

OpenStudy (anonymous):

when you say simultaneous do you mean set them equal to each other and solve for x?

OpenStudy (anonymous):

the slope would be 2 because we treat the (1,1) as the 0 point right?

hartnn (hartnn):

oh yes, we put x=1 in 1-3/x^2 to get slope =1-3/1 =1-3 = -2 so slope is actually -2 get this ?

hartnn (hartnn):

and yes your equation y=-2x+3 is correct, now i said to equate the two curve equation to get the intersection points so -2x+3 = x+3/x can u find values of x from here ?

OpenStudy (anonymous):

x=0 and x=-1 so which is the correct answer?

hartnn (hartnn):

x is not 0 -2x^2 +3x=x^2+3 so 3x^2 -3x + 3 =0 or x^2-x+1 =0 solve this quadratic equation to get the values of x

OpenStudy (anonymous):

x=1

hartnn (hartnn):

that equation doesn't seem to have real roots. can u please verify your question ?

OpenStudy (anonymous):

This is the question. I hit the exact same wall but when I gave DNE as the answer it was incorrect and now I just don't know what to do! :(

hartnn (hartnn):

the line -2x+3 and curve x+3/x do not intersect.

OpenStudy (anonymous):

so this question is impossible? They wouldn't give me an impossible question Grrrr

hartnn (hartnn):

i will try to go through this again, if i come up with something else, i will let you know

OpenStudy (anonymous):

Thanks a bunch!